# Robust Point Matching with Distance Profiles

### v2

@@ 2 Outlier Robustness of Distance Profile Matching
$W_{1}\left(\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\right)\leq\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}.$
Hence, we have
$|W(i,j)-W^{\prime}(i,j)|\leq\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}+\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}=:U+V.$
Fix $i\in[n]$ such that $t(i)\in[K]$. Then,
$\begin{split}\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W(i,j)&\geq\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W^{\prime}(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W^{\prime}(i,j)-2(U+V)\\ &\geq\min_{\beta\in[s]\backslash\{t(i)\}}\overline{W}(t(i),\beta)-\overline{W}(t(i),t(i))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\Omega-2R(Z_{1}+Z_{2})-2(U+V).\end{split}$
Therefore, $2R(Z_{1}+Z_{2})+2(U+V)<\Omega$ implies that $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. We show that $\mathbb{P}(E)\leq\delta$, where $E$ is the event $(2R(Z_{1}+Z_{2})+2(U+V)\geq\Omega)$. By Lemma 1 and (2.3), we have
$\mathbb{P}\left(\sum_{k=1}^{t}\left|\frac{n_{k}}{n}-p_{k}\right|\geq\frac{\Omega}{8R}\right),\mathbb{P}\left(\sum_{k=1}^{s}\left|\frac{m_{k}}{m}-q_{k}\right|\geq\frac{\Omega}{8R}\right)\leq\frac{\delta}{4}.$
Since $\xi_{i}-\xi_{\ell}\,|\,t(i),t(\ell)\sim N(0,(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})I_{d})$ and thus $\|\xi_{i}-\xi_{\ell}\|_{2}^{2}\,|\,t(i),t(\ell)\sim(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})\chi^{2}(d)$ for $i\neq\ell$, one can deduce that
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Omega^{2}}{128\sigma^{2}}\right).$
By Lemma 2 and (2.4), we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Omega^{2}}{128\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{\Omega^{2}}{2048\sigma^{2}}\right)\leq\frac{\delta}{4}.$
Similarly, we have
$\mathbb{P}\left(\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\frac{\delta}{4}.$
@@
$\mathbb{P}(E)\leq\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{8R}\right)+\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{8R}\right)+\mathbb{P}\left(U\geq\frac{\Omega}{8}\right)+\mathbb{P}\left(V\geq\frac{\Omega}{8}\right)\leq\delta.$
Hence, we have (2.1). ∎
@@
So far, we have focused on the matching of the points from the common part, namely, the first $K$ components. The matching guarantee ensures that all $X_{i}$’s such that $t(i)\in[K]$ are correctly matched, while ignoring the points from the outlier components. In this model, if we assume a stronger condition than (2.2), namely,
$\left(\min_{\alpha\neq\beta}\overline{W}(\alpha,\beta)\wedge\min_{\alpha\in[t]\backslash[K],\beta\in[s]}\overline{W}(\alpha,\beta)\right)-\max_{\alpha\in[K]}\overline{W}(\alpha,\alpha)=:\Omega^{\prime}>0,$ (2.5)
and replace $\Omega$ in (2.3), (2.4) with $\Omega^{\prime}$, we can show that
$\max_{i\in[n]:t(i)\in[K]}W(i,\pi(i))<\min_{i\in[n]:t(i)\notin[K]}W(i,\pi(i))$ (2.6)
@@ 3 Noise Stability of Assignment-Based Matching
For the case where there is no rigid transformation, namely, $T$ is the identity map, the noise stability guarantee in the above model is well-studied in the literature, known as the feature matching problem. In this case, the simplest approach is to solve the discrete optimal transport problem [Peyré and Cuturi, 2019] using the squared Euclidean distances, namely,

### v5

@@ 2 Outlier Robustness of Distance Profile Matching
$W_{1}\bigg{(}\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}},\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|\eta_{s(j)}-\eta_{s(\ell)}\|_{2}}\bigg{)}\leq\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}=:V.$
Hence, $|W(i,j)-W^{\prime}(i,j)|\leq U+V$ and $|W(i,j)-\overline{W}(t(i),s(j))|\leq R(Z_{1}+Z_{2})+U+V$. Now, fix $i\in[n]$ such that $t(i)\in[K]$. Then,
$\begin{split}&\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}W(i,j)-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}W(i,j)\\ &\geq\min_{\begin{subarray}{c}j\in[m]\\ s(j)\neq t(i)\end{subarray}}\overline{W}(t(i),s(j))-\min_{\begin{subarray}{c}j\in[m]\\ s(j)=t(i)\end{subarray}}\overline{W}(t(i),s(j))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\min_{\beta\in[s]\backslash\{t(i)\}}\overline{W}(t(i),\beta)-\overline{W}(t(i),t(i))-2R(Z_{1}+Z_{2})-2(U+V)\\ &\geq\underbrace{\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}}_{=:\Omega}-2R(Z_{1}+Z_{2})-2(U+V),\end{split}$ (2.7)
where the last inequality follows from Lemma 1. Therefore, $2R(Z_{1}+Z_{2})+2(U+V)<\Omega$ implies that $s(\pi(i))=t(i)$ for any $i\in[n]$ such that $t(i)\in[K]$. We show that $\mathbb{P}(E)\leq\delta$, where $E$ is the event $(2R(Z_{1}+Z_{2})+2(U+V)\geq\Omega)$. By Lemma 3, we have
$\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{8R}\right)$ $\leq 2K(t-1)\exp\left(-\frac{2n\Omega^{2}}{(8R)^{2}}\right),$
$\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{8R}\right)$ $\leq 2K(s-1)\exp\left(-\frac{2m\Omega^{2}}{(8R)^{2}}\right).$
Meanwhile, for $i\neq\ell$, note that we have $\xi_{i}-\xi_{\ell}\,|\,t(i),t(\ell)\sim N(0,(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})I_{d})$, which implies $\|\xi_{i}-\xi_{\ell}\|_{2}^{2}\,|\,t(i),t(\ell)\sim(\sigma_{t(i)}^{2}+\sigma_{t(\ell)}^{2})\chi^{2}(d)$. Hence, we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{(\Omega/8)^{2}}{2\sigma^{2}}\right).$
As $\Omega^{2}\geq 256d\cdot\max_{k\in[t]}\sigma_{k}^{2}$, we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i}-\xi_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{(\Omega/8)^{2}}{2\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\sigma^{2}}\right).$
Similarly, as $\Omega^{2}\geq 256d\cdot\max_{k\in[s]}\tau_{k}^{2}$, we have
$\mathbb{P}\left(\max_{j,\ell\in[m]}\|\zeta_{j}-\zeta_{\ell}\|_{2}\geq\frac{\Omega}{8}\right)\leq m^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\tau^{2}}\right).$
@@
$\begin{split}\mathbb{P}(E)&\leq\mathbb{P}\left(Z_{1}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(Z_{2}\geq\frac{\Omega}{4R}\right)+\mathbb{P}\left(U\geq\frac{\Omega}{8}\right)+\mathbb{P}\left(V\geq\frac{\Omega}{8}\right)\\ &\leq 2K(t+s-2)\exp\left(-\frac{2(n\wedge m)\Omega^{2}}{(8R)^{2}}\right)+2(n\vee m)^{2}\exp\left(-\frac{(\Omega/8)^{2}}{32\Gamma^{2}}\right)\\ &\leq\frac{\delta}{2}+\frac{\delta}{2}=\delta,\end{split}$
where the last inequality follows from (2.3). Hence, we have (2.2). ∎
@@
From the proof of Theorem 1, one may have noticed that we can replace the left-hand side of (2.3) with
$\Omega_{o}:=\min_{\alpha\in[K]}\left(\min_{\beta\in[s]\backslash\{\alpha\}}\overline{W}(\alpha,\beta)-\overline{W}(\alpha,\alpha)\right).$
To see this, it suffices to observe that we can replace $\Omega$ in the last line of (2.7) with $\Omega_{o}$, and the rest of the proof follows the same with $\Omega_{o}$ in place of $\Omega$. Since $\Omega_{o}\geq\Omega$, the condition (2.3) is a stronger condition than requiring $\Omega_{o}$ to be at least the right-hand side of (2.3). As we briefly mentioned before the proof, we chose to present (2.3) as it is written in terms of the outlier proportion which is more interpretable.
Lastly, we provide the lemmas used in the proof of Theorem 1.
Lemma 1.
For the centers $\theta_{1},\ldots,\theta_{t}$ and $\eta_{1},\ldots,\eta_{s}$ in Assumption 1, define $\overline{W}$ as in (2.4). Then, for any $k\in[K]$, we have
$\overline{W}(\alpha,\alpha)\leq\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)}\cdot\max_{(\alpha,\beta)\in[t]\times[s]}\|\theta_{\alpha}-\eta_{\beta}\|_{2}.$ (2.8)
Proof.
The Wasserstein-1 distance on $\mathscr{P}(\mathbb{R})$ can be written as the $L^{1}$ distance between cumulative distribution functions. For $\alpha\in[K]$, we have
$\begin{split}\overline{W}(\alpha,\alpha)&=\int_{\mathbb{R}}\left|\sum_{k=1}^{t}p_{k}1\{\|\theta_{\alpha}-\theta_{k}\|_{2}\leq z\}-\sum_{k=1}^{s}q_{k}1\{\|\eta_{\alpha}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &=\int_{\mathbb{R}}\left|\sum_{k\in[t]\backslash[K]}p_{k}1\{\|\theta_{\alpha}-\theta_{k}\|_{2}\leq z\}-\sum_{k\in[s]\backslash[K]}q_{k}1\{\|\theta_{\alpha}-\eta_{k}\|_{2}\leq z\}\right|\,\mathrm{d}z\\ &=(1-\lambda)W_{1}\left(\sum_{k\in[t]\backslash[K]}\tfrac{p_{k}}{1-\lambda}\delta_{\|\theta_{\alpha}-\theta_{k}\|_{2}},\sum_{k\in[s]\backslash[K]}\tfrac{q_{k}}{1-\lambda}\delta_{\|\theta_{\alpha}-\eta_{k}\|_{2}}\right),\end{split}$
where $\lambda=\sum_{k\in[K]}p_{k}$. Since $|\|\theta_{\alpha}-\theta_{k}\|_{2}-\|\theta_{\alpha}-\eta_{\ell}\|_{2}|\leq\|\theta_{k}-\eta_{\ell}\|_{2}$ for any $k\in[t]\backslash[K]$ and $\ell\in[s]\backslash[K]$, one can deduce that (2.8) holds. ∎
Lemma 2.
Let $z_{1}\leq\cdots\leq z_{t}$ be any real numbers. For any $(p_{1},\ldots,p_{t}),(q_{1},\ldots,q_{t})\in\Delta_{t}$, we have
$W_{1}\bigg{(}\sum_{k=1}^{t}p_{k}\delta_{z_{k}},\sum_{k=1}^{t}q_{k}\delta_{z_{k}}\bigg{)}\leq(z_{t}-z_{1})\cdot\max_{j\in[t-1]}\bigg{|}\sum_{k=1}^{j}(p_{k}-q_{k})\bigg{|}.$ (2.9)
Proof.
Notice that
$W_{1}\bigg{(}\sum_{k=1}^{t}p_{k}\delta_{z_{k}},\sum_{k=1}^{t}q_{k}\delta_{z_{k}}\bigg{)}=\int_{\mathbb{R}}\bigg{|}\sum_{k=1}^{t}(p_{k}-q_{k})1\{z_{k}\leq z\}\bigg{|}\,\mathrm{d}z=\sum_{j=1}^{t-1}\bigg{|}\sum_{k=1}^{j}(p_{k}-q_{k})\bigg{|}\cdot(z_{j+1}-z_{j}),$
where the last equality follows from
$\bigg{|}\sum_{k=1}^{t}(p_{k}-q_{k})1\{z_{k}\leq z\}\bigg{|}=\sum_{j=1}^{t-1}\bigg{|}\sum_{k=1}^{j}(p_{k}-q_{k})\bigg{|}\cdot 1\{z_{j}\leq z<z_{j+1}\}\quad\forall z\in\mathbb{R}.$
Therefore, we have (2.9). ∎
Necessary Noise Level for the Matching Guarantee
In Theorem 1, one can intuitively notice that the noise level, denoted as $\Gamma=\max_{k\in[t]}\sigma_{k}\vee\max_{k\in[s]}\tau_{k}$, must be small enough to ensure the matching guarantee. While its desirable scale in relation to the sample size is transparent from the analysis, one may be curious about its magnitude compared to the ambient dimension $d$. From the first term on the right-hand side of (2.3), one can see that $\Gamma$ has to be of order $d^{-1/2}$. At first, such a dimension-dependence may look like a strong condition, it is in fact a minimal condition given that both sides of (2.3) scale with typical Euclidean distances in $\mathbb{R}^{d}$. See Figure 2. Each Gaussian component essentially corresponds to a ball of radius $\Gamma\sqrt{d}$, which has to be well-separated from the other balls to ensure that it is a distinct component. Accordingly, it is natural to have $\Gamma<\frac{\|c_{1}-c_{2}\|_{2}}{\sqrt{d}}$ for any distinct $c_{1},c_{2}\in\{\theta_{1},\ldots,\theta_{t},\eta_{1},\ldots,\eta_{s}\}$. This explains that $\Gamma=O(d^{-1/2})$ implied by (2.3) is a minimal condition for the matching guarantee.
Figure 2: Illustration of the Gaussian mixtures. Each component is represented as a ball centered at the corresponding center, where the radius of the ball is roughly $\Gamma\sqrt{d}$. All of them are contained in a larger ball of diameter $R$.
Remark 2.
So far, we have focused on the matching of the points from the common part, namely, the first $K$ components. The matching guarantee ensures that all $X_{i}$’s such that $t(i)\in[K]$ are correctly matched, while ignoring the points from the outlier components. If we assume a stronger condition by replacing the left-hand side of (2.3) with the following quantity
$\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)\wedge\min_{\begin{subarray}{c}\alpha\in[t]\backslash[K]\\ \beta\in[s]\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)},$ (2.10)
we can show that
$\max_{\begin{subarray}{c}i\in[n]\\ t(i)\in[K]\end{subarray}}W(i,\pi(i))<\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}W(i,\pi(i))$ (2.11)
@@ 3 Noise Stability of Assignment-Based Matching
For the case where there is no rigid transformation, namely, $T$ is the identity map, the noise stability guarantee in the above model is well-studied in the literature, known as the feature matching problem. In this case, the simplest approach is to solve the discrete optimal transport problem [30] using the squared Euclidean distances, namely,